OpenStudy (anonymous):

calculate the density of a rectangular piece of wood from the following data: Length: 25cm Width: 10cm Height: 4cm Mass: 15g calculate the density of a rectangular piece of wood from the following data: Length: 25cm Width: 10cm Height: 4cm Mass: 15g @Physics

OpenStudy (anonymous):

m/v so 15g/4cm x 10cm x 25 cm 15g/1000cm^3 0.015g/cm^3

OpenStudy (ashish56):

first, find volume = 25*10*4=1000 Now, density = 15/1000=0.015